# Simultaneously search for multi-target Galactic binary gravitational waves in reduced parameter space with LMPSO-CV

### v1

@@ Abstract
We propose an innovative approach to the concurrent exploration of gravitational waves originating from Galactic binaries through the development of a new Local Maxima Particle Swarm Optimization (LMPSO) algorithm. Our methodology employs strategic Create Voids (CV) to streamline parameter space, maximizing the identification of local maxima for the $\mathcal{F}$-statistic even in the overlapped signals case. Subsequently, a “$-\mathcal{F}$-statistic-analysis”, which implements the astrophysical models and properties of $\mathcal{F}$-statistic in parameter space, is conducted to reveal Galactic binary gravitational wave signals within the dataset. Our new approach eliminates inaccuracies associated with signal subtraction contamination, which is a challenge for traditional iterative-subtraction methods when addressing low signal-to-noise ratio signals (e.g., SNR $<$ 15). To demonstrate the efficacy of our approach, we utilize the residuals from the LISA mock data challenge (LDC1-4), where 10982 injection sources with optimal SNR $>$ 15 have been eliminated. The LMPSO-CV method efficiently identifies 8995 signals with a 47.7% false source fraction or 3463 signals with a 26.9% false source fraction when the correlation coefficient threshold is set to 0.8.
@@ I Introduction
In the evolving landscape of ground-based gravitational wave (GW) detectors, recent observations [1, 2, 3] underscore advancements in sensitivity and increased detection distances, resulting in a surge of GW events [4, 5]. However, the inherent frequency limitations (10 Hz to 1000 Hz) of these detectors introduce a bias toward specific sources, such as stellar-mass binary black hole mergers, binary neutron star mergers, and neutron star-black hole mergers [6, 7, 8]. Despite their significance, these detections merely scratch the surface of the comprehensive GW spectrum. The quest for detecting GWs at alternative frequencies has prompted the exploration of kilohertz GWs [9, 10], lunar GW detection [11], space-based GW detection [12, 13], and the utilization of pulsar timing arrays [14, 15, 16]. Pulsar timing arrays, in particular, have yielded compelling evidence of potential signals in the extremely low-frequency range [17, 18, 19]. Beyond pulsar timing arrays, several space-based GW detection initiatives, including LISA [12, 13, 20], Taiji [21, 22, 23, 24], and Tianqin [25, 26, 27, 28], among others, are in preparation, expanding the horizons of gravitational wave research. The Taiji program [21, 22, 23, 24] and the LISA program [12, 13, 20] are positioned to establish space-based GW detectors strategically, with one located 20 degrees ahead and the other 20 degrees behind Earth’s orbit. Configured with arm lengths of approximately 3 million kilometers and 2.5 million kilometers, respectively, each detector consists of three independent satellites interconnected by laser links. Operating on the same fundamental principles as ground-based GW detectors, these instruments will measure minute spatial displacements between pairs of satellites for GW detection. By evading ground-based seismic noise and possessing substantial arm lengths, space-based GW detectors exhibit sensitivities several orders of magnitude lower than their ground-based counterparts, covering a frequency range from about $10^{-4}$ Hz to $10^{-1}$ Hz. Within this frequency band, an array of GW sources is anticipated, including supermassive black hole mergers, Galactic binaries, and extreme mass ratio inspirals, among others [29]. Galactic binaries, especially double white dwarf (DWD) binaries, are expected to be abundant. According to current Galactic models, tens of millions to billions of Galactic binaries could emit nearly monochromatic GW signals within the sensitivity frequencies of LISA and Taiji [30]. The prevalence of these signals poses a significant challenge in signal resolution due to potential overlap. Additionally, as Galactic binary gravitational wave signals persist stably throughout the detection period, they have the potential to impact the analysis of other types of signals. In recent decades, many data processing methods have emerged for analyzing GW signals from Galactic binaries [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. Predominantly grounded in the $\mathcal{F}$-statistic, this approach significantly reduces the parameters requiring exploration for monochromatic gravitational waves, with some techniques directly probing all parameters defining these waves. The majority of existing methodologies adopt an iterative-subtraction strategy when dealing with multiple signals. This strategy entails identifying a signal, subtracting it, and then searching for subsequent signals in the residual. The iterative-subtraction process persists until no signal above the threshold is detected. However, as the number of identified signals increases and their signal-to-noise ratio (SNR) diminishes, this strategy becomes susceptible to introducing inaccuracies in the subtraction process, thereby contaminating the residual data [42]. This contamination complicates the search for signals with low SNR. Our objective is to identify signals with a SNR lower than 15 with respect to instrument noise. While the iterative-subtraction strategy proves effective in scenarios where signals have an SNR exceeding 15, benefiting from a relatively sparse signal density and high SNR, its efficacy diminishes for signals with SNR below 15. In cases where numerous signals coexist amid a high density, the iterative-subtraction strategy encounters challenges. To address this issue, we propose a novel search strategy that eliminates reliance on iterative-subtraction, implementing it on LDC1-4 data. Prior to initiating the search, assuming a method exists to eliminate signals with SNR $>$ 15, we start by subtracting all injected Galactic binary signals with optimal SNR $>$ 15 from the raw data. Subsequently, we proceed with our method within the residual. Our algorithm is designed to identify as many local maxima of the $\mathcal{F}$-statistic surpassing a specified threshold as possible in parameter space. Following this, we analyze the results and extract signal parameters from these local maxima. The structure of this article is organized as follows: In Sec. II, we provide an introduction to the gravitational waveform of DWD binaries and the response function of the LISA detector, both based on the LDC1-4 data. Sec. III explores the distribution of the $\mathcal{F}$-statistic when single or multiple signals coexist in LISA data, leading to the generation of degeneracy noise. Sec. IV outlines the LMPSO-CV method utilized for searching the local maxima of the $\mathcal{F}$-statistic, incorporating a specially developed Local Maxima Particle Swarm Optimization (LMPSO) algorithm and a technique for reducing parameter space to prevent redundant searches for the same position in parameter space—referred to as “create voids” (CV). In Sec. V, we elaborate on how to analyze these local maxima with the “$-\mathcal{F}$-statistic-analysis”, focusing on removing a significant amount of degeneracy noise and mitigating instrument noise. Sec. VI presents the results in terms of the false source fraction and SNR, along with the parameter errors and computational time. Conclusions are drawn in Sec. VII. In Appendix A, we present the results of the case search.
@@ II.1 Galactic binaries distribution
where $\mathcal{A}$ represents the GW amplitude, $\iota$ represents the Galactic binary orbit to the line of sight from the Solar System barycentric (SSB) origin, $\phi_{0}$ represents the initial phase. $f$ is the GW frequency, and $\dot{f}$ is the first order derivative of $f$. Incorporating the polarization, denoted as $\psi$, from $h_{+}$ and $h_{\times}$, along with two positional parameters—Ecliptic Latitude $\beta$ and Ecliptic Longitude $\lambda$—constitutes a total of eight parameters that characterize a Galactic binary system. There are 29857650 Galactic binaries in LDC1-4 data, but only 27680 with optimal SNR $>$ 7 in two-year detection. Among these, there are 10982 binaries with optimal SNR $>$ 15, which are sources that most of them can already solved good by current search methods. For the remaining 16698 sources, detection is more difficult because they are mainly concentrated in low ($f<1\times 10^{-3}$ Hz) and medium ($1\times 10^{-3}$ Hz $<f<$ $4\times 10^{-3}$ Hz) frequencies with a high density and are more susceptible to instrument noise.
Figure 1: Galactic binaries distribution in the galactic coordinate system of LDC1-4 (optimal SNR $>$ 7). The color bar represents the number of binaries within one square degree on the celestial sphere. The two red lines mark the area where the Galactic Latitude is between -0.5 rad and 0.5 rad ($-28.65\degree\sim 28.65\degree$), and there are 27,436 binaries in this area out of 27,680 binaries in total. This prior information will be used in the data analysis (See Sec. V.2.3 for details).
Figure 2: Distribution of Galactic binaries in different frequency bins of the LDC1-4 data. Each bin has a width of 0.01 mHz. We categorize the signals into two groups based on optimal SNR: those with SNR $>$ 15 and those with 7 $<$ SNR $<$ 15. In the figure, we observe that binaries are more concentrated at low frequencies ($f<1\times 10^{-3}$ Hz) and medium frequencies ($1\times 10^{-3}$ Hz $<$ $f$ $<$ $4\times 10^{-3}$ Hz) when considering signals with 7 $<$ optimal SNR $<$ 15.
Fig. 1 illustrates the distribution of binaries in the galactic coordinate system, with the majority congregating on the Galactic Disk. Less than one percent of binaries exhibit a Galactic Latitude outside the range of -0.5 rad to 0.5 rad. In Fig. 2, we depict the frequency distribution of binaries. Most binaries with SNR $>$ 15 are readily resolvable using current search methods. In contrast, binaries with 7 $<$ SNR $<$ 15 exhibit higher densities at low frequencies ($f<1\times 10^{-3}$ Hz) and medium frequencies ($1\times 10^{-3}$ Hz $<$ $f$ $<$ $4\times 10^{-3}$ Hz), posing challenges for effective discrimination. Given our primary focus on signals with low SNR (SNR $<$ 15), we initiate the process by subtracting binary waveforms with optimal SNR $>$ 15 from the raw data.
@@ II.2 Space-based GW detector response
The development of a comprehensive response for a space-based gravitational wave detector is a complex task, requiring consideration of intricate details related to the detectors’ orbits. In this regard, we rely on the formulation presented in Ref. [36], a model that has demonstrated high effectiveness in previous studies [36, 38, 42]. While the gravitational wave (GW) emitted by a Galactic binary inherently exhibits a monochromatic waveform, the Doppler effect introduces frequency broadening due to the orbital motion of the detector. The Earth, moving at a velocity of approximately 30 km/s around the SSB, shares a similar speed with LISA. As expressed by the Doppler effect formula, $f_{1}=f\times\frac{c\pm v}{c}$ (where c is the speed of light), the GW frequency undergoes a shift of approximately $0.01\%$ in one frequency direction (this is a maximum, and the specific shift depends on the location of the GW source).
@@ III.1 $\mathcal{F}$-statistic
where $\omega$ is the angular GW frequency and $L$ is the arm length of LISA. $a^{(k)}=\left(a^{(1)},a^{(2)},a^{(3)},a^{(4)}\right)$ is the reparametrization of four extrinsic parameters $\mathcal{A}$, $\phi_{0}$, $\psi$ and $\iota$. $h^{(k)}(t)$ is a matrix of template waveforms and it depends on the four remaining intrinsic parameters: $f$, $\dot{f}$, $\beta$ and $\lambda$. The $\mathcal{F}$-statistic is proposed to search for monochromatic GWs [48]. Supposing there is a Galactic binary GW in the data, the detector’s data in $X$ can be expressed as
@@
The symbol $\tilde{}$ represents the Fourier transform, $*{}^{*}$ denotes complex conjugation, and $S_{h}$ represents the one-sided spectral density of the detector’s noise. It maximizes the log-likelihood function $\ln\Lambda$ concerning parameters $a^{(k)}$, by solving
@@
III.2 $\mathcal{F}$-statistic degeneracy in the parameter space
Figure 3: The x-y coordinate plane represents the ecliptic plane, with the x-axis directed toward the vernal equinox. LISA orbits the SSB within the ecliptic plane at a speed denoted by $\vec{\nu}$, and the origin $o$ corresponds to the SSB’s position. Within the LISA data, there exists a GW signal originating from binary $A$ with parameters $f_{A}$, $\beta_{A}$ and $\lambda_{A}$. Employing a match filtering method, such as the $\mathcal{F}$-statistic, for detection involves assuming that the matched waveform corresponds to binary $A^{\prime}$ ($f_{A^{\prime}},\leavevmode\nobreak\ \beta_{A^{\prime}},\leavevmode\nobreak\ \lambda_{A^{\prime}}$).
Due to the inherent nature of the $\mathcal{F}$-statistic as a matching filtering method, employing a LISA response GW waveform to correlate with the signal in LISA data yields a $\mathcal{F}$-statistic value that exceeds instrument noise levels if the two waveforms share the same Doppler modulating frequency at the same time. In Fig. 3, the x-y coordinate plane denotes the ecliptic plane, with the x-axis directed toward the vernal equinox. LISA orbits the SSB with a velocity vector $\vec{\nu}$ within the ecliptic plane (this is a simplified model and does not include cartwheel motions of LISA), and the SSB is positioned at the origin ($o$). Assuming binary $A$ is the signal in the LISA data, $f_{A}$ represents binary $A$’s frequency relative to the SSB. Due to the orbital motion of LISA, the frequency $f_{A}$, modulated by the Doppler effect $f_{A_{do}}$ in the LISA data, can be expressed as
$f_{A_{do}}=f_{A}\frac{c+\nu_{A}}{c},$ (10)

### v3

@@ Abstract
The search for Galactic binary gravitational waves is a critical challenge for future space-based gravitational wave detectors, such as LISA. We propose an innovative approach to simultaneously explore gravitational waves originating from Galactic binaries by developing a new Local Maxima Particle Swarm Optimization (LMPSO) algorithm. Our method identifies local maxima in the $\mathcal{F}$-statistic and applies astrophysical models and the properties of the $\mathcal{F}$-statistic within parameter space to uncover Galactic binary gravitational wave signals in the dataset. This new approach effectively addresses the inaccuracies often associated with signal subtraction contamination, a challenge for traditional iterative subtraction methods, particularly when dealing with low signal-to-noise ratio (SNR) signals (e.g., SNR $<$ 15). We also account for the effects of overlapping signals and degeneracy noise. To demonstrate the effectiveness of our approach, we use residuals from the LISA mock data challenge (LDC1-4), where 10,982 injected sources with SNR $\geq$ 15 have been removed. For the remaining low-SNR sources (SNR $<$ 15), our method efficiently identifies 6,508 signals, achieving a 63.2% detection rate.
@@ I Introduction
In the evolving landscape of ground-based gravitational wave (GW) detectors, recent observations [1, 2, 3] underscore advancements in sensitivity and increased detection distances, resulting in a surge of GW events [4, 5]. However, the inherent frequency limitations (10 Hz to 1000 Hz) of these detectors introduce a bias toward specific sources, such as stellar-mass binary black hole mergers, binary neutron star mergers, and neutron star-black hole mergers [6, 7, 8]. Despite their significance, these detections merely scratch the surface of the comprehensive GW spectrum. The quest for detecting GWs at alternative frequencies has prompted the exploration of kilohertz GWs [9, 10], lunar GW detection [11], space-based GW detection [12, 13], and the utilization of pulsar timing arrays [14, 15, 16]. Pulsar timing arrays, in particular, have yielded compelling evidence of potential signals in the extremely low-frequency range [17, 18, 19]. Beyond pulsar timing arrays, several space-based GW detection initiatives, including LISA [12, 13, 20], Taiji [21, 22, 23, 24], and Tianqin [25, 26, 27, 28], among others, are in preparation, expanding the horizons of gravitational wave research.
The Taiji program [21, 22, 23, 24] and the LISA program [12, 13, 20] are positioned to establish space-based GW detectors strategically, with one located 20 degrees ahead and the other 20 degrees behind Earth’s orbit. Configured with arm lengths of approximately 3 million kilometers and 2.5 million kilometers, respectively, each detector consists of three independent satellites interconnected by laser links. Operating on the same fundamental principles as ground-based GW detectors, these instruments will measure minute spatial displacements between pairs of satellites for GW detection. By evading ground-based seismic noise and possessing substantial arm lengths, space-based GW detectors exhibit sensitivities several orders of magnitude lower than their ground-based counterparts, covering a frequency range from about $10^{-4}$ Hz to $10^{-1}$ Hz. Within this frequency band, an array of GW sources is anticipated, including supermassive black hole mergers, Galactic binaries, and extreme mass ratio inspirals, among others [29]. Galactic binaries, especially double white dwarf (DWD) binaries, are expected to be abundant. According to current Galactic models, tens of millions to billions of Galactic binaries could emit nearly monochromatic GW signals within the sensitivity frequencies of LISA and Taiji [30]. The prevalence of these signals poses a significant challenge in signal resolution due to potential overlap. Additionally, as Galactic binary gravitational wave signals persist stably throughout the detection period, they have the potential to impact the analysis of other types of signals.
In recent decades, many data processing methods have emerged for analyzing GW signals from Galactic binaries [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 39, 44]. Among these algorithms, LDC1-4 data (Radler) were utilized for testing in three different algorithms. The first algorithm employed RJMCMC’s global fit approach [37, 43], while the second utilized MCMC and GPU-accelerated techniques [39, 44], and the third implemented a PSO-based iterative-subtraction algorithm combined with a cross-validation approach [38, 42]. When applied to the LDC1-4 data, the first algorithm identified 826, 1819, 4356, and 7255 sources over periods of 1.5, 3, 6, and 12 months, respectively [43]. The second algorithm detected 3407, 6251, and 10363 sources over periods of 6, 12, and 24 months, respectively [44]. The third algorithm identified 10341 sources over 24 months [38]. Additionally, the performance of the third algorithm was also evaluated on 3 and 6 months data, revealing the detection of 3010 and 5418 sources (this evaluation was conducted without cross-validation, which could improve the detection rate but lead to the loss of some sources) [42]. Despite the efforts of these algorithms to target sources with SNR $>7$ or SNR $>10$, there are still many sources that remain unsearched, given that there are approximately 27,000 injection sources with SNR $>7$ and 18,000 with SNR $>10$ for two-year detection. This scenario may arise due to several factors. Firstly, the presence of high SNR sources may overshadow the weaker low SNR signals during multi-source fitting, resulting in their neglect. Secondly, as the number of searched signals increases, the potential for inaccurate subtraction contamination from the iterative-subtraction algorithm in the data residual grows, impacting the detection of low SNR signals. Thirdly, the more serious overlap of low SNR signals poses a more significant challenge compared to high SNR signals, particularly in the absence of sufficient priors regarding the number of sources and the approximate parameter ranges. To alleviate these challenges, especially the second, we propose a non iterative-subtraction strategy and evaluate its performance in single-source fitting scenarios. We believe that our method holds advantages in low SNR and low-frequency regions characterized by higher source densities compared to high SNR and high-frequency regions.
Our objective is to accurately identify signals with SNR $<$ 15 (with respect to instrument noise) during the two-year detection period. While the iterative-subtraction strategy proves effective in scenarios where signals have SNR $\geq$ 15 [38, 42], benefiting from relatively sparse signal density and high SNR, its efficacy gradually diminishes as the SNR decreases. We propose a novel search strategy capable of simultaneously detecting multiple Galactic binaries, and implement it using LDC1-4 data. Before initiating the search, assuming a method exists to eliminate signals with SNR $\geq$ 15, we begin by subtracting all injection sources with SNR $\geq$ 15 from the LDC1-4 data. Subsequently, we proceed with our method within the residual data. Our algorithm is designed to identify as many local maxima of the $\mathcal{F}$-statistic surpassing a specified threshold as possible in parameter space. Following this, we analyze the results and extract signal parameters from these local maxima.
The structure of this article is organized as follows: In Sec. II, we provide an introduction to the gravitational waveform of DWD binaries and the response function of the LISA detector, both based on the LDC1-4 data. Sec. III explores the distribution of the $\mathcal{F}$-statistic when single or multiple signals coexist in LISA data, which leads to the generation of degeneracy noise, and discusses the presence of overlapping signals within the LDC1-4 injection sources. Sec. IV outlines the LMPSO-CV method utilized for searching the local maxima of the $\mathcal{F}$-statistic, incorporating a specially developed Local Maxima Particle Swarm Optimization (LMPSO) algorithm and a technique for reducing parameter space to prevent redundant searches for the same position in parameter space—referred to as “create voids” (CV). In Sec. V, we elaborate on how to analyze these local maxima with the “$-\mathcal{F}$-statistic-analysis”, focusing on removing a significant amount of degeneracy noise and mitigating instrument noise. Sec. VI presents the results in terms of the detection rate and SNR, along with the parameter errors and computational time. Conclusions are drawn in Sec. VII. In Appendix A, we present the results of the case search.
@@ II.1 Galactic binaries distribution
where $\mathcal{A}$ represents the GW amplitude, $\iota$ represents the Galactic binary orbit to the line of sight from the Solar System barycentric (SSB) origin, $\phi_{0}$ represents the initial phase. $f$ is the GW frequency, and $\dot{f}$ is the first order derivative of $f$. Incorporating the polarization, denoted as $\psi$, from $h_{+}$ and $h_{\times}$, along with two positional parameters—Ecliptic Latitude $\beta$ and Ecliptic Longitude $\lambda$—constitutes a total of eight parameters that characterize a Galactic binary system.
There are 29857650 Galactic binaries in LDC1-4 data, but only 27680 with SNR $>$ 7 in two-year detection. Among these, there are 10982 binaries with SNR $>$ 15, which are sources that most of them can already solved good by current search methods. For the remaining 16698 sources, detection is more difficult because they are mainly concentrated in low ($f<1\times 10^{-3}$ Hz) and medium ($1\times 10^{-3}$ Hz $<f<$ $4\times 10^{-3}$ Hz) frequencies with a high density and are more susceptible to instrument noise.
Figure 1: Galactic binaries distribution in the galactic coordinate system of LDC1-4 (SNR $>$ 7). The color bar represents the number of binaries within one square degree on the celestial sphere. The two red lines mark the area where the Galactic Latitude is between -0.5 rad and 0.5 rad ($-28.65\degree\sim 28.65\degree$), and there are 27,436 binaries in this area out of 27,680 binaries in total. This prior information will be used in the data analysis (See Sec. V.2.3 for details).
Figure 2: Distribution of Galactic binaries in different frequency bins of the LDC1-4 data. Each bin has a width of 0.01 mHz. We categorize the signals into two groups based on SNR: those with SNR $>$ 15 and those with 7 $<$ SNR $<$ 15. In the figure, we observe that binaries are more concentrated at low frequencies ($f<1\times 10^{-3}$ Hz) and medium frequencies ($1\times 10^{-3}$ Hz $<$ $f$ $<$ $4\times 10^{-3}$ Hz) when considering signals with 7 $<$ SNR $<$ 15.
Fig. 1 illustrates the distribution of binaries in the galactic coordinate system, with the majority congregating on the Galactic Disk. Less than one percent of binaries exhibit a Galactic Latitude outside the range of -0.5 rad to 0.5 rad. In Fig. 2, we depict the frequency distribution of binaries. Most binaries with SNR $>$ 15 are readily resolvable using current search methods. In contrast, binaries with 7 $<$ SNR $<$ 15 exhibit higher densities at low frequencies ($f<1\times 10^{-3}$ Hz) and medium frequencies ($1\times 10^{-3}$ Hz $<$ $f$ $<$ $4\times 10^{-3}$ Hz), posing challenges for effective discrimination. Given our primary focus on signals with low SNR (SNR $<$ 15), we initiate the process by subtracting binary waveforms with SNR $>$ 15 from the raw data.
@@ II.2 Space-based GW detector response
The development of a comprehensive response for a space-based gravitational wave detector is a complex task, requiring consideration of intricate details related to the detectors’ orbits. In this regard, we rely on the formulation presented in Ref. [36], a model that has demonstrated high effectiveness in previous studies [36, 38, 42].
While the gravitational wave (GW) emitted by a Galactic binary inherently exhibits a monochromatic waveform, the Doppler effect introduces frequency broadening due to the orbital motion of the detector. The Earth, moving at a velocity of approximately 30 km/s around the SSB, shares a similar speed with LISA. As expressed by the Doppler effect formula, $f_{1}=f\times\frac{c\pm v}{c}$ (where c is the speed of light), the GW frequency undergoes a shift of approximately $0.01\%$ in one frequency direction (this is a maximum, and the specific shift depends on the location of the GW source).
@@ III.1 $\mathcal{F}$-statistic
where $\omega$ is the angular GW frequency and $L$ is the arm length of LISA. $a^{(k)}=\left(a^{(1)},a^{(2)},a^{(3)},a^{(4)}\right)$ is the reparametrization of four extrinsic parameters $\mathcal{A}$, $\phi_{0}$, $\psi$ and $\iota$. $h^{(k)}(t)$ is a matrix of template waveforms and it depends on the four remaining intrinsic parameters: $f$, $\dot{f}$, $\beta$ and $\lambda$.
The $\mathcal{F}$-statistic is proposed to search for monochromatic GWs [50]. Supposing there is a Galactic binary GW in the data, the detector’s data in $X$ can be expressed as
@@
The symbol $\tilde{}$ represents the Fourier transform, denotes complex conjugation, and $S_{h}$ represents the one-sided spectral density of the detector’s noise.
It maximizes the log-likelihood function $\ln\Lambda$ concerning parameters $a^{(k)}$, by solving
@@
III.2 The frequency-position parameter degeneracy
Figure 3: The $x-y$ coordinate plane represents the ecliptic plane, with the $x$-axis directed toward the vernal equinox. LISA orbits the SSB within the ecliptic plane at a speed denoted by $\vec{\nu}$, and the origin $o$ corresponds to the SSB’s position. Within the LISA data, there exists a GW signal originating from binary $A$ with parameters $f_{A}$, $\beta_{A}$ and $\lambda_{A}$. Employing a match filtering method, for detection involves assuming that the match waveform corresponds to binary $A^{\prime}$ ($f_{A^{\prime}},\leavevmode\nobreak\ \beta_{A^{\prime}},\leavevmode\nobreak\ \lambda_{A^{\prime}}$).
Due to the inherent nature of the $\mathcal{F}$-statistic as a matching filtering method, employing a LISA response GW waveform to correlate with the signal in LISA data yields a $\mathcal{F}$-statistic value that exceeds instrument noise levels if the two waveforms share the same Doppler modulating frequency at the same time. This results in the frequency-position parameter degeneracy. In Fig. 3, the $x-y$ coordinate plane denotes the ecliptic plane, with the $x$-axis directed toward the vernal equinox. LISA orbits the SSB with a velocity vector $\vec{\nu}$ within the ecliptic plane (this is a simplified model and does not include cartwheel motions of LISA), and the SSB is positioned at the origin ($o$). Assuming binary $A$ is the signal in the LISA data, $f_{A}$ represents binary $A$’s frequency relative to the SSB. Due to the orbital motion of LISA, the frequency $f_{A}$, modulated by the Doppler effect $f_{A_{\text{do}}}$ in the LISA data, can be expressed as
$f_{A_{\text{do}}}=f_{A}\frac{c+\nu_{A}}{c},$ (10)